OpenStudy (anonymous):

I want to verify if my answer is right for this integral: ∫_0^(1)(2x-1)^20 dx is this 0.047619?

OpenStudy (anonymous):

are you raising zero to a power??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

= (1/42) ( 1 - ( -1) ) = 1/21

OpenStudy (anonymous):

right that the same as my answer then I divided it so is 0.047619